# Identifying Majorana edge and end modes in Josephson junction of $p$-wave superconductor with magnetic barrier

### v1

@@ IV.1 For $p_{x}+p_{y}$ type pairing
First, we consider the $p_{x}+p_{y}$-type pairing and illustrate the behaviour of $N(\mathbf{r},0)$ in the $x-y$ plane as depicted in Fig. 5. In Fig. 5(a) and 5(b), we plot the $N(\mathbf{r},0)$ in the hybrid topological phase for $\phi=\pi/2$ and $\phi=\pi$, respectively fixing $\mu=3t$, $t_{m}=2t,$ and $\xi_{m}=5$. Interestingly, in the hybrid topological phase the MZMs that are localized at the ends of the magnetic barrier, are now gapped out due to the applied phase bias as shown in Fig. 5(a). This suggests that tuning the phase difference, the system moves from hybrid topological phase to the phase with only MFEMs. Similar to $\phi=0$ [Fig. 2(c)], at $\phi=\pi/2$, the junction MFEMs still remains gapped. Then, in Fig. 5(b) for $\phi=\pi$, we also observe the absence of MZMs at the ends of magnetic barrier. However, now a bulk channel appears near the junction which is localized along the $y$-direction. Making analogy with 1D $p$-wave SCs, we find the energy splitting of junction localized Majorana modes, $\Delta E_{jun}\sim\cos(\phi/2)$ Alicea (2012). Therefore, for $\phi=\pi$, the junction MFEMs becomes degenerate at zero energy and appears to be localized along the junction irrespective of the barrier strength, as can be seen in the LDOS profile of Fig. 5(b). In contrast to the hybrid topological phase, in Fig. 5(c), when only MZMs exist, the LDOS exhibits no significant difference compared to $\phi=0$ case [see Fig. 2 (d)]. Moreover, a channel forms through the magnetic barrier as the band gap closes at $\phi=\pi$ as shown in Fig. 5(d). This effectively connects the localized end modes through the bulk. This feature is observed in both Figs. 5(b) and 5(d). Additionally, in Fig. 5(c) and (d), the MFEMs are completely gapped out.
@@ IV.2 For $p_{x}+ip_{y}$ type pairing
Here, we discuss the effect of phase bias on LDOS profile considering $p_{x}+ip_{y}$-type SC and illustrate our findings in Fig. 6. In Fig. 6(a) and 6(b), we show zero energy LDOS in the $x-y$ plane for $\phi=\pi/2$ and $\phi=\pi$ respectively with $\mu=t,t_{m}=2t,$ and $\xi_{m}=5$. In Fig. 6(a), we observe the presence of MZMs at the top end of the magnetic barrier along with the MEMs localized along the edges of the system. Interestingly, for the same values of $\mu,t_{m}$, and $\xi_{m}$ the system hosts only MEMs in absence of any phase difference i.e.,  $\phi=0$. Therefore, for $p_{x}+ip_{y}$-type pairing, introducing a phase bias drives the system into hyrid topological phase where both MEMs and MZMs coexist. Next, for $\phi=\pi$, the system hosts MEMs at the junction along the $y$-direction as depicted in Fig.6(b). This has the similar origin as the $p_{x}+p_{y}$-type pairing where the junction MEMs becomes gapless for $\phi=\pi$. At $\phi=\pi/2$, Fig. 6(c) exhibits similar behavior to the $p_{x}+p_{y}$-type pairing [see Fig 5(c)] where the topological region corresponds to only MZMs. While at $\phi=\pi$, the intermediate channel coexists with MZMs in the absence of edge modes [see Fig. 6(d)]. The reason is at $\phi=\pi$, the spectrum becomes gapless and this channel appears due to the formation of junction localized bulk states.
@@ V.2 Effect of Magnetic Barrier Strength
VI Summary and discussion
@@
Acknowledgments

### v2

@@ III.2 For $p_{x}+ip_{y}$ type pairing
In this case, $p_{x}+ip_{y}$ pairing breaks time reversal symmetry and we compute the Chern number using the Fukui method Fukui et al. (2005) to topologically characterize the phase. Specifically, we consider the case where only MEMs are present, with no spin-dependent tunneling present (i.e., $t_{m}=0$). As discussed in Appendix B.1, the Chern number takes the value $C=1$ in the region $0<\mu<4t$, which corresponds to the edge localized dispersive MEMs depicted in Fig. 3(a).
@@ IV.1 For $p_{x}+p_{y}$ type pairing
First, we consider the $p_{x}+p_{y}$-type pairing and illustrate the behaviour of $N(\mathbf{r},0)$ in the xyx-y plane as depicted in Fig. 5. In Fig. 5(a) and 5(b), we plot the $N(\mathbf{r},0)$ in the hybrid topological phase for $\phi=\pi/2$ and $\phi=\pi$, respectively fixing $\mu=3t$, $t_{m}=2t,$ and $\xi_{m}=5$. Interestingly, in the hybrid topological phase the MZMs that are localized at the ends of the magnetic barrier, are now gapped out due to the applied phase bias as shown in Fig. 5(a). This suggests that tuning the phase difference, the system moves from hybrid topological phase to the phase with only MFEMs. Similar to $\phi=0$ [Fig. 2(c)], at $\phi=\pi/2$, the junction MFEMs still remains gapped. Then, in Fig. 5(b) for $\phi=\pi$, we also observe the absence of MZMs at the ends of magnetic barrier. However, now a bulk channel appears near the junction which is localized along the $y$-direction. Making analogy with 1D $p$-wave SCs, we find the energy splitting of junction localized Majorana modes, $\Delta E_{jun}\sim\cos(\phi/2)$ Alicea (2012). Therefore, for $\phi=\pi$, the junction MFEMs becomes degenerate at zero energy and appears to be localized along the junction irrespective of the barrier strength, as can be seen in the LDOS profile of Fig. 5(b). In contrast to the hybrid topological phase, in Fig. 5(c), when only MZMs exist, the LDOS exhibits no significant difference compared to $\phi=0$ case [see Fig. 2 (d)]. Moreover, a channel forms through the magnetic barrier as the band gap closes at $\phi=\pi$ as shown in Fig. 5(d). This effectively connects the localized end modes through the bulk. This feature is observed in both Figs. 5(b) and 5(d). Additionally, in Fig. 5(c) and (d), the MFEMs are completely gapped out.
@@ IV.2 For $p_{x}+ip_{y}$ type pairing
Here, we discuss the effect of phase bias on LDOS profile considering $p_{x}+ip_{y}$-type SC and illustrate our findings in Fig. 6. In Fig. 6(a) and 6(b), we show zero energy LDOS in the xyx-y plane for $\phi=\pi/2$ and $\phi=\pi$ respectively with $\mu=t,t_{m}=2t,$ and $\xi_{m}=5$. In Fig. 6(a), we observe the presence of MZMs at the top end of the magnetic barrier along with the MEMs localized along the edges of the system. Interestingly, for the same values of $\mu,t_{m}$, and $\xi_{m}$ the system hosts only MEMs in absence of any phase difference i.e.,  $\phi=0$. Therefore, for $p_{x}+ip_{y}$-type pairing, introducing a phase bias drives the system into hyrid topological phase where both MEMs and MZMs coexist. Next, for $\phi=\pi$, the system hosts MEMs at the junction along the $y$-direction as depicted in Fig.6(b). This has the similar origin as the $p_{x}+p_{y}$-type pairing where the junction MEMs becomes gapless for $\phi=\pi$. At $\phi=\pi/2$, Fig. 6(c) exhibits similar behavior to the $p_{x}+p_{y}$-type pairing [see Fig 5(c)] where the topological region corresponds to only MZMs. While at $\phi=\pi$, the intermediate channel coexists with MZMs in the absence of edge modes [see Fig. 6(d)]. The reason is at $\phi=\pi$, the spectrum becomes gapless and this channel appears due to the formation of junction localized bulk states.
@@ V.2 Effect of Magnetic Barrier Strength
In the case of the $p_{x}+p_{y}$-wave superconductor, we observe a $4\pi$-periodic Josephson effect at $t_{m}=0$, which diminishes as $t_{m}$ increases. This behavior is consistent with the 1D scenario involving two $p$-wave Kitaev chains coupled via a magnetic weak link as illustrated in Appendix C. The jump at $\phi=\pi$ and $t_{m}=0$ becomes less sharp due to the finite-size effects in 2D. In contrast, for the $p_{x}+ip_{y}$-wave superconductor, we do not observe a $4\pi$ periodicity even at $t_{m}=0$. The reason can be attributed to the fact that in a Josephson junction between two 2D chiral superconductors of the same chirality, both the energy spectrum and the JC exhibit a $2\pi$ periodicity Kwon et al. (2004). Furthermore, in both $p_{x}+p_{y}$ and $p_{x}+ip_{y}$ type superconductors, higher harmonics appear in the current-phase relation (CPR). This behavior aligns with the expectations as spin-dependent tunneling, which effectively introduces a Zeeman field along the $z$-direction (see Appendix A). The presence of such effective magnetic field at the junction can lead to deviations from the conventional sinusoidal CPR and gives rise to higher harmonic contributions in the JC Willsch et al. (2024); Fukaya et al. (2022); Goldobin et al. (2007).
VI Stability of MZMs and Josephson Current
In this section, we investigate the stability of MZMs in a $p_{x}+ip_{y}$-wave SC by introducing a localized perturbation at the end points of the magnetic barrier. This perturbation modifies the tunneling Hamiltonian $H_{LR}$, as defined in Eq. (3), and is applied specifically at the barrier terminations located at $y=0\text{ and }y=L_{y},\text{ with }x=L_{x}/2$.
The corresponding perturbative Hamiltonian is given by:
[…]
where $\tilde{t}_{1}$ denotes the strength of the perturbation.
We consider a strong perturbation with $\tilde{t}_{1}=10t$. Our numerical results exhibit that the MZMs remain robust in the presence of such local perturbation. As shown in Fig. 9(a), the MZMs are not destroyed. However, they are slightly displaced from the ends of the magnetic barrier to adjacent lattice sites along the junction. Nevertheless, the characteristic zero-energy feature associated with MZMs continues to appear in the LDOS, indicating their persistence. In Fig. 9(b), we depict the JC for both the unperturbed ($\tilde{t}_{1}=0$) and perturbed ($\tilde{t}_{1}=10t$) cases. The CPR remains qualitatively unchanged, further confirming the robustness of the hump signature indicating the mixed phase (presence of both MEMs and MZMs) for $p_{x}+ip_{y}$ SC.
Figure 9: (a) In panel (a), we showcase the persistence of MZMs in LDOS spectra under a perturbation of strength $\tilde{t}_{1}=10t$. In panel (b), we show the JC for two cases: $\tilde{t}_{1}=0$ (red curve) and $\tilde{t}_{1}=10t$ (green curve). We use the other model parameters as $\mu=2t$, $t_{m}=2t$, $t_{0}=t$, $\Delta_{0}=0.5t$, $\xi_{m}=5t$, and $t=1$, on a finite size system of $50\times 50$ lattice sites.
VII Summary and discussion
@@ Acknowledgments
Appendix A Derivation of the Effective 1D Hamiltonian
In this appendix, we derive an effective 1D Hamiltonian starting from Eq. (1). To simplify the problem, we consider a system with $L_{x}=2$ and $L_{y}\rightarrow\infty$, applying periodic boundary condition (PBC) along the $y$-direction to approximate the bulk limit. This setup reduces the system to two coupled superconducting chains aligned along the $y$-axis with magnetic weak link between them. We aim to develop a simplified 1D model that can be solved exactly to provide analytical insights into our main results.
The system consists of left ($L$) and right ($R$) superconducting chains coupled via a harmonic spin-textured barrier $V_{LR}(y)$. The $L$ and $R$ regions are characterized by superconducting phases $\phi_{L}=-\phi/2$ and $\phi_{R}=+\phi/2$, respectively. The total Hamiltonian can then be expressed as Sardinero et al. (2024)
$H=\begin{pmatrix}H_{L}&V_{LR}(y)\\ V_{LR}^{\text{\textdagger}}(y)&H_{R}\end{pmatrix}\$, (11)
where, the left and right Hamiltonians $H_{L,R}$ are given by
[…]
$-\sin\phi_{L,R}\tau_{y}\sigma_{y}$]\ ,
and the inter-chain coupling reads
$V_{LR}=-t_{0}\tau_{z}\sigma_{0}-t_{m}\left(\sin\theta_{y}\,\tau_{0}\sigma_{x}+\cos\theta_{y}\,\tau_{0}\sigma_{z}\right)\ .$
Combining these, the full Hamiltonian becomes
[…]
$+\sin\frac{\phi}{2}\,u_{z}\tau_{y}\sigma_{y}]-t_{0}u_{x}\tau_{z}\sigma_{0}-t_{m}(\sin\theta_{y}\,u_{x}\tau_{0}\sigma_{x}$
$+\cos\theta_{y}\,u_{1}\tau_{0}\sigma_{3})\ ,$
where $u$ represents Pauli matrices acting on the left and right subspace.
To remove the $y$-dependence from the angle $\theta_{y}$, we perform a local unitary transformation using the operator
[…]
The effective Hamiltonian in the rotated frame is then given by $H_{\text{eff}}=U^{\text{\textdagger}}HU$. For the low-energy, we approximate:
$\cos k_{y}\approx 1-\partial_{y}^{2},\quad\sin k_{y}\approx-i\partial_{y}\ ,$
Substituting these into the transformed Hamiltonian yields
$H_{\text{eff}}$ =($=-\left[\mu_{\text{eff}}+2t\cos k_{y}\right]u_{0}\tau_{z}\sigma_{0}-\frac{2\pi t}{\xi_{m}}\sin k_{y}\,u_{0}\tau_{z}\sigma_{y}$ (14)
$\quad+\Delta\cos\frac{\phi}{2}\left(\sin k_{y}u_{0}\tau_{x}\sigma_{y}+\frac{\pi}{\xi_{m}}u_{0}\tau_{x}\sigma_{0}\right)-t_{0}u_{x}\tau_{z}\sigma_{0}$
$\quad-t_{m}u_{x}\tau_{0}\sigma_{z}+\Delta\sin\frac{\phi}{2}\left(\sin k_{y}u_{z}\tau_{y}\sigma_{y}+\frac{\pi}{\xi_{m}}u_{z}\tau_{y}\sigma_{0}\right)\ .$
Note that, the effective chemical potential is given by $\mu_{\text{eff}}=\mu-\frac{\pi^{2}t}{\xi_{m}^{2}}$, indicating a reduction by the term $\frac{\pi^{2}t}{\xi_{m}^{2}}$. Effective spin-orbit coupling emerges in the Hamiltonian, $H_{\text{eff}}$ through the term $\frac{2\pi t}{\xi_{m}}\sin k_{y}$. Additionally, the superconducting pairing is modified by an effective factor of $\frac{\pi}{\xi_{m}}$. An effective Zeeman term also arises from spin dependent hopping strength, $t_{m}$, along the $z$ direction.
Appendix B Topological Characterization
In this appendix, we present the topological invariants associated with MEMs and MZMs, characterized respectively by the Chern number and the winding number.
B.1 Chern Number
We compute the Chern number using the Fukui method Fukui et al. (2005) in the absence of spin-dependent tunneling ($t_{m}=0$), where only the normal tunneling amplitude ($t_{0}$) is present. In this regime, the system behaves as a single 2D $p$-wave superconductor hosting MEMs localized at its boundary. For the $p_{x}+ip_{y}$-type pairing, the Chern number is calculated and shown in Fig. 10. In contrast, for the $p_{x}+p_{y}$ configuration, the bulk spectrum becomes gapless, rendering the Chern number ill-defined.
For the $p_{x}+ip_{y}$ SC, the Chern number takes the value,
  • $C=0$ in the trivial regime,
  • $C=+1$ for $0<\mu<4t$, corresponding to the topological regime hosting a single chiral MEM, and
  • C=1C=-1 in another topological phase depending on the chemical potential,
as illustrated in Fig. 10.
Figure 10: Chern number is depicted as a function of chemical potential $\mu$ (along the range of $\mu$ mentioned in Fig. 3(a) of the main text) for the $p_{x}+ip_{y}$-wave SC when $t_{m}=0$.
B.2 Winding Number
To characterize the MZMs, we analyze the effective 1D model (Eq. (14)) and compute the winding number. For simplicity, we set the superconducting phase $\phi=0$. We define a chiral symmetry operator $C=u_{0}\tau_{y}\sigma_{y}$, under which the Hamiltonian can be recast into an off-diagonal form via a unitary transformation $U_{c}$, composed of the eigenvectors of $C$. In the chiral basis, the Hamiltonian takes the form
$H=\begin{pmatrix}0&H^{+}(k)\\ H^{-}(k)&0\end{pmatrix}\$, (15)
where, $H(k)$ denotes the $4\times 4$ matrices defined within the chiral subspace.
The winding number $W$, a topological invariant, is then calculated as Chiu et al. (2016); Mondal et al. (2023b):
[…]
We evaluate $W$ in the plane of the spin-dependent tunneling $t_{m}$ and the spatial modulation vector $\xi_{m}$, for two values of the chemical potential: $\mu=3t$ and $\mu=4.2t$ (see Figs. 11(a) and (b) respectively). These correspond to hybrid and fully topological phases in the full 2D lattice model, respectively. The computed winding number takes values $W=1$ or $2$, indicating the presence of one or two MZMs at each end of the effective 1D junction along $y$ direction.
Figure 11: We illustrate the variation of winding number $W$ in $t_{m}-\xi_{m}$ plane for the effective 1D model choosing $\mu=3t$ in panel (a) and in panel (b) considering $\mu=4.2t$. The other model parameters are considered to be $\Delta=0.5t$, $t_{0}=t$ and $t=1$.
Appendix C Magnetic weak link between two 1D $p$-wave SCs
In this appendix, we numerically study the behavior of a Josephson junction comprised of a magnetic weak link connecting two 1D $p$-wave Kitaev chains as shown in Fig. 12. Our focus is on how varying the magnetic coupling strength, denoted by $t_{m}$, influences the energy spectrum and the corresponding JC.
Initially, when $t_{m}=0$, the system hosts localized MZMs at the junction (with energy $E_{\rm jun}$) and outer ends (with energy $E_{\rm end}$). When the phase difference becomes $\phi=\pi$, two branches of $E_{\rm jun}$ crosses each other, signaling the change in fermionic parity of the ground state resulting in the characteristic $4\pi$-periodic energy-phase relation as illustrated in Fig. 13(a).
Figure 12: We schematically depict the Josephson junction comprised of two 1D $p$-wave Kitaev chains with superconducting phase difference of $\phi$, connected via a magnetic weak link with strength $t_{m}$.
Figure 13: We display the energy-phase ($E$-$\phi$) relation displaying the evolution of energy levels with phase difference $\phi$. Panel (a) $t_{m}=0$: protected zero-energy crossing at $\phi=\pi$ is depicted. In panels (b), (c), (d) and (e) we consider $t_{m}\neq 0$: gap opens, lifting the degeneracy. In panel (f), we set $t_{m}=t_{0}$ and there is no phase effect.
Figure 14: We showcase the JC ($I(\phi)$) as a function of phase bias ($\phi$) considering the 1D $p$-wave $J$ setup in presence of a magnetic link with various strengths $t_{m}$.
When the magnetic coupling $t_{m}$ is turned on, the zero-energy crossing at $\phi=\pi$ disappears as illustrated in Figs. 13(b)-(f). This indicates the hybridization of the MZMs localized at the junction, lifting the degeneracy and resulting in a gap opening at $\phi=\pi$, thereby destroying the $4\pi$-periodicity in the energy spectrum.
This transition is also reflected in the behavior of JC. As shown in Fig. 14, when $t_{m}=0$, the current manifests a $4\pi$-periodic component (discontinous jump) due to the presence of topologically protected MZMs. However, even a small $t_{m}\ll t_{0}$ suppresses this behavior, restoring the conventional $2\pi$-periodicity. As the magnetic coupling approaches the normal tunneling strength ($t_{m}\approx t_{0}$), the JC eventually becomes vanishingly small (see Fig. 14).